# Benchmarking M-LTSF: Frequency and Noise-Based Evaluation of Multivariate Long Time Series Forecasting Models

### v1

@@ Abstract
Understanding the robustness of deep learning models for multivariate long-term time series forecasting (M-LTSF) remains challenging, as evaluations typically rely on real-world datasets with unknown noise properties. We propose a simulation-based evaluation framework that generates parameterizable synthetic datasets, where each dataset instance corresponds to a different configuration of signal components, noise types, signal-to-noise ratios, and frequency characteristics. These configurable components aim to model real-world multivariate time series data without the ambiguity of unknown noise. This framework enables fine-grained, systematic evaluation of M-LTSF models under controlled and diverse scenarios. We benchmark four representative architectures S-Mamba (state-space), iTransformer (transformer-based), R-Linear (linear), and Autoformer (decomposition-based). Our analysis reveals that all models degrade severely when lookback windows cannot capture complete periods of seasonal patters in the data. S-Mamba and Autoformer perform best on sawtooth patterns, while R-Linear and iTransformer favor sinusoidal signals. White and Brownian noise universally degrade performance with lower signal-to-noise ratio while S-Mamba shows specific trend-noise and iTransformer shows seasonal-noise vulnerability. Further spectral analysis shows that S-Mamba and iTransformer achieve superior frequency reconstruction. This controlled approach, based on our synthetic and principle-driven testbed, offers deeper insights into model-specific strengths and limitations through the aggregation of MSE scores and provides concrete guidance for model selection based on signal characteristics and noise conditions.
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
I Introduction
@@ II Related Work
II-A Long Time Series Forecasting Architectures
@@
State Space Models: Another emerging approach to time series forecasting involves modeling the underlying system using a state space model. The core model proposed by Gu et al. (S4[20] efficiently handles variable-length sequences while maintaining linear complexity with respect to sequence length thus giving them an advantage over Transformers. Expanding on this foundation, the Mamba architecture of Gu et al. [21] introduces a novel input-dependent, time-varying state space model that enables efficient processing of long sequences. Unlike traditional models, Mamba dynamically adapts its state transitions based on the input, allowing it to capture complex temporal dependencies. Its built-in selective mechanism may facilitate attention to relevant historical information while filtering out irrelevant or noisy data. Building upon this foundation, the S-Mamba architecture [22] adapts Mamba’s selective state space mechanism specifically for time series forecasting by replacing the attention computation in the iTransformer architecture with bidirectional Mamba blocks. This substitution maintains the pattern recognition capabilities while achieving near-linear computational complexity, making it particularly suitable for efficient multivariate time series prediction.
II-B Benchmark Datasets
Recent research on M-LTSF methods is commonly based on a set of standard benchmark datasets for evaluation. The ETT (Electricity Transformer Temperature) dataset presented by Zhou et al.  [17] contains 2-year electricity transformer data from China with variants for hourly and 15-minute sampling rates. Other commonly used datasets include electricity consumption data from 321 consumers, daily exchange rates from eight countries, hourly traffic occupancy rates from California freeways, and meteorological data with 21 indicators recorded every 10 minutes  [18, 39].
@@ II-B Benchmark Datasets
II-C Noise analysis and modeling in M-LTSF
A critical challenge in M-LTSF is the tendency of deep learning methods to overfit on noise present in the training data, particularly when the signal-to-noise ratio is low [36]. Traditional approaches to noise analysis in time series forecasting rely on oversimplified assumptions. Singh [40] evaluated noise impact by only injecting Gaussian noise and filtering based on a cutoff frequency that classified low-frequency content as ”trend” energy and high frequencies as noise. This approach is problematic because real-world noise is rarely only Gaussian in nature, and the binary classification of frequency components oversimplifies the complex spectral characteristics of time series, where meaningful patterns can exist across the entire frequency spectrum. Further, recent works on probabilistic forecasting [41] and noise robustness [42, 13] have shown that models vary in their ability to maintain performance under noisy or adversarial conditions. Wang et al.[43] demonstrated the value of synthetic data for controlled evaluation of Transformer architectures across ten signal types, though their analysis was limited to Transformer-based models and combined multiple noise types simultaneously rather than isolating individual noise characteristics.
III M-LTSF Benchmark Framework
@@ III-B Noise Components
III-B1 Signal Independent Noise
@@
III-B2 Signal Dependent Noise
@@
III-C Multivariate Time Series Generation
@@ IV Experimental Design
IV-A Evaluation Models
For our evaluation, we selected four architecturally diverse M-LTSF models based on their distinct design principles and recent performance in the literature. These models include S-Mamba as a state-space model, iTransformer as an attention-based approach, R-Linear as a linear method, and Autoformer utilizing a decomposition-based design (see Table I). The models differ in both size and training efficiency: Autoformer is the largest with 14.6 Million parameters but has the fastest training time of 75 seconds per epoch. Attempts to increase Autoformer’s parameters further resulted in out-of-memory errors on our smaller compute resources. In comparison, R-Linear has 7.5M parameters and requires 225 seconds per epoch, while S-Mamba and iTransformer have 9.2M and 6.4M parameters, with training times of 219 and 216 seconds per epoch, respectively. The hyperparameters for each model, as well as those used in the experiments, are available in the associated GitHub repository.
TABLE I: Overview of evaluated M-LTSF Models
@@ IV-A Evaluation Models
Autoformer [17] Correlation , running avg. 14.6M 278.2 514.6
iTransformer [19] Attention (, instance norm.) 6.4M 96.6 325.4
R-Linear [24] Linear (, instance norm.) 7.5M 92.7 1001.0
S-Mamba [22] SSM (, instance norm.) 9.2M 95.3 270.7
IV-B Evaluation Approach
@@ IV-C Parameter Space
IV-D Dataset Instance Generation
@@ V Evaluation
V-A Noise and Trend-Free Baseline Evaluation
@@
V-B Trend Component Impact
@@
To get a better understanding how trend components in the signal influence model performances, trend components were introduced to the synthetic data for a smaller subset of frequencies (frequency ranges: 1000-1500, 8000-8500, 16000-16500) to balance meticulousness and computational effort. A consistent improvement in model performance across all architectures compared to purely seasonal signals (see Figure 8) could be observed. This improvement can possibly be attributed to the directional guidance that trend components provide during gradient-based optimization. Purely seasonal signals can create challenging optimization landscapes due to their inherent symmetries and the phase prediction requirements that arise when models must learn to match periodic patterns precisely. The trend component serves as an ”anchor” that helps models establish temporal direction and reduces the phase prediction burden that is particularly challenging in purely periodic signals. The relative performance rankings between models remain consistent with trend addition, indicating that the fundamental signal type preferences are architectural characteristics rather than artifacts of the no-trend condition.
V-C Noise Robustness Analysis

### v2

@@ Abstract
Understanding the robustness of deep learning models for multivariate long-term time series forecasting (M-LTSF) remains challenging, as evaluations typically rely on real-world datasets with unknown noise properties. We propose a simulation-based evaluation framework that generates configurable synthetic dataset instances based on controllable signal components, noise types, signal-to-noise ratios, and frequency characteristics. These configurable components aim to model real-world multivariate time series data without the ambiguity of unknown noise. This framework enables fine-grained, systematic evaluation of M-LTSF models under controlled and diverse scenarios. We benchmark four representative architectures S-Mamba, iTransformer, R-Linear, and Autoformer. Our analysis shows that model performance deteriorates sharply when lookback windows fail to capture full seasonal cycles. S-Mamba and Autoformer perform best on sawtooth patterns, whereas R-Linear and iTransformer prefer sinusoidal signals. White and Brownian noise consistently impair performance at lower signal-to-noise ratios. Additionally, S-Mamba is sensitive to trend noise, and iTransformer to seasonal noise. Spectral analysis further indicates that S-Mamba and iTransformer offer superior frequency reconstruction. This controlled approach, based on our synthetic evaluation framework, offers deeper insights into model-specific strengths and limitations through the aggregation of MSE scores and provides concrete guidance for model selection based on signal characteristics and noise conditions.
Keywords Multivariate Long Time Series Forecasting (M-LTSF) $\cdot$ Time series analysis $\cdot$ Signal Processing $\cdot$ Noise Analysis $\cdot$ Spectral Analysis
1 Introduction
@@ 2 Related Work
2.1 Long Time Series Forecasting Architectures
@@
State Space Models: Another emerging approach to time series forecasting involves modeling the underlying system using a state space model. The core model proposed by Gu et al. [16] (S4) efficiently handles variable-length sequences while maintaining linear complexity with respect to sequence length thus giving them an advantage over Transformers. Expanding on this foundation, the Mamba architecture of Gu et al. [15] introduces a novel input-dependent, time-varying state space model that enables efficient processing of long sequences. Unlike traditional models, Mamba dynamically adapts its state transitions based on the input, allowing it to capture complex temporal dependencies. Its built-in selective mechanism may facilitate attention to relevant historical information while filtering out irrelevant or noisy data. Building upon this foundation, the S-Mamba architecture [39] adapts Mamba’s selective state space mechanism specifically for time series forecasting by replacing the attention computation in the iTransformer architecture with bidirectional Mamba blocks. This substitution maintains the pattern recognition capabilities while achieving near-linear computational complexity, making it particularly suitable for efficient multivariate time series prediction.
2.2 Benchmark Datasets
Recent research on M-LTSF methods is commonly based on a set of standard benchmark datasets for evaluation. The ETT (Electricity Transformer Temperature) dataset presented by Wu et al. [41] contains 2-year electricity transformer data from China with variants for hourly and 15-minute sampling rates. Other commonly used datasets include electricity consumption data from 321 consumers, daily exchange rates from eight countries, hourly traffic occupancy rates from California freeways, and meteorological data with 21 indicators recorded every 10 minutes  [45, 20].
@@ 2.2 Benchmark Datasets
2.3 Noise analysis and modeling in M-LTSF
A critical challenge in M-LTSF is the tendency of deep learning methods to overfit on noise present in the training data, particularly when the signal-to-noise ratio is low [22]. Traditional approaches to noise analysis in time series forecasting rely on oversimplified assumptions. Singh [34] evaluated noise impact by only injecting Gaussian noise and filtering based on a cutoff frequency that classified low-frequency content as "trend" energy and high frequencies as noise. This approach is problematic because real-world noise is rarely only Gaussian in nature, and the binary classification of frequency components oversimplifies the complex spectral characteristics of time series, where meaningful patterns can exist across the entire frequency spectrum. Further, recent works on probabilistic forecasting [42] and noise robustness [9, 7] have shown that models vary in their ability to maintain performance under noisy or adversarial conditions. Wang et al. [38] demonstrated the value of synthetic data for controlled evaluation of Transformer architectures across ten signal types, though their analysis was limited to Transformer-based models and combined multiple noise types simultaneously rather than isolating individual noise characteristics.
3 M-LTSF Benchmark Framework
@@ 3.2 Noise Components
3.2.1 Signal Independent Noise
@@
3.2.2 Signal Dependent Noise
@@
3.3 Multivariate Time Series Generation
@@ 4 Experimental Design
4.1 Evaluation Models
For our evaluation, we selected four architecturally diverse M-LTSF models based on their distinct design principles and recent performance in the literature. These models include S-Mamba as a state-space model, iTransformer as an attention-based approach, R-Linear as a linear method, and Autoformer utilizing a decomposition-based design (see Table 1). The models differ in both size and training efficiency: Autoformer is the largest with 14.6 Million parameters but has the fastest training time of 75 seconds per epoch. Attempts to increase Autoformer’s parameters further resulted in out-of-memory errors on our smaller compute resources. In comparison, R-Linear has 7.5M parameters and requires 225 seconds per epoch, while S-Mamba and iTransformer have 9.2M and 6.4M parameters, with training times of 219 and 216 seconds per epoch, respectively. The hyperparameters for each model, as well as those used in the experiments, are available in the associated GitHub repository.
Table 1: Overview of evaluated M-LTSF Models
@@ 4.1 Evaluation Models
Autoformer Correlation , running avg. 14.6M 278.2 514.6
iTransformer Attention (, instance norm.) 6.4M 96.6 325.4
R-Linear Linear (, instance norm.) 7.5M 92.7 1001.0
S-Mamba SSM (, instance norm.) 9.2M 95.3 270.7
4.2 Evaluation Approach
@@ 4.3 Parameter Space
4.4 Dataset Instance Generation
@@ 5 Evaluation
5.1 Noise and Trend-Free Baseline Evaluation
@@
5.2 Trend Component Impact
@@
To get a better understanding how trend components in the signal influence model performances, trend components were introduced to the synthetic data for a smaller subset of frequencies (frequency ranges: 1000-1500, 8000-8500, 16000-16500) to balance meticulousness and computational effort. A consistent improvement in model performance across all architectures compared to purely seasonal signals (see Figure 8) could be observed. This improvement can possibly be attributed to the directional guidance that trend components provide during gradient-based optimization. Purely seasonal signals can create challenging optimization landscapes due to their inherent symmetries and the phase prediction requirements that arise when models must learn to match periodic patterns precisely. The trend component serves as an "anchor" that helps models establish temporal direction and reduces the phase prediction burden that is particularly challenging in purely periodic signals. The relative performance rankings between models remain consistent with trend addition, indicating that the fundamental signal type preferences are architectural characteristics rather than artifacts of the no-trend condition.
5.3 Noise Robustness Analysis